OpenStudy (anonymous):

The pilot of an airplane traveling 190 km/h wants to drop supplies to flood victims isolated on a patch of land 100 m below. How many seconds before the plane is directly overhead should the supplies be dropped?

OpenStudy (anonymous):

100 m .... s=1/2*a*t^2 100 = 1/2 * 9.81*t^2 t^2 = 200/9.81 t = sqrt(200/9.81) We have assumed no air drag.

OpenStudy (anonymous):

And no parachute for the box either... it'll just slam into the ground... ;P